OpenStudy (anonymous):

The perimeter of a rectangle is fifty centimeters. The length of the rectangle is twenty centimeters. What is the width of the rectangle?

OpenStudy (jdoe0001):

the perimeter of a rectangle is the length + length + width + width = 50 we know that the length is 25, so replace each length with 25 and solve for "width"

OpenStudy (anonymous):

so 25+25= 50

OpenStudy (anonymous):

i mean 20+20= 40

OpenStudy (jdoe0001):

hmm sorry I was caught up on something

OpenStudy (jdoe0001):

length = l width = w thus perimeter will be l + l + w + w = 50 but we know that l = 20, thus 20 + 20 + w + w = 50 40 + 2w = 50 # solve for "w"

OpenStudy (anonymous):

can you please just give me the answer this is the only question i have left on my test

OpenStudy (jdoe0001):

I could but i don't see how that "helps" you think about it this way, what if the teacher hands you out the test along with the answers too what's the point of the test at all?

OpenStudy (jdoe0001):

the idea behind exercises is not to trancribe material, so much it's to understand it

OpenStudy (anonymous):

true :/ ah ok i'll try to solve it

OpenStudy (anonymous):

20 + 20 + 5 + 5 = 50 so the answer is 5?

OpenStudy (jdoe0001):

20 + 20 + w + w = 50 40 + 2w = 50 what if you subtract 40 from each side, the left and right side, what would that give you?